OpenStudy (anonymous):

how do you find the sum of the infinite geometric series: (5/6)+(1/6)+(1/30)+...

OpenStudy (anonymous):

first term is \(\frac{5}{6}\) and common ratio is \(\frac{1}{5}\) use \[\sum_{k=0}^{\infty}ar^k=\frac{a}{1-r}\] with \(a=\frac{5}{6}\) and \(r=\frac{1}{5}\)